OpenStudy (anonymous):

Calculate the molarity of the solution. 29 g of NaCl (g.f.w. = 58.5 g/mol) in 0.030 liters. 0.015 M 17 M 67 M

OpenStudy (anonymous):

First you want to find the moles of 29g of NaCl. Then you can plug it into the equation for molarity which is moles of solute/liter of solution.

OpenStudy (anonymous):

how do you do that???

OpenStudy (anonymous):

Find the moles of NaCl?

OpenStudy (anonymous):

How do you do that?!?

OpenStudy (anonymous):

29 g of Nacl * molar mass of NaCl= answer

OpenStudy (anonymous):

I think she needs you to set it up in a equation

OpenStudy (anonymous):

I just did.

OpenStudy (anonymous):

Then you take that answer and put it over the liters of solution.

OpenStudy (anonymous):

this is my question.... how would you find the molar mass of NaCl?

OpenStudy (anonymous):

The molar mass is the sum of the atomic mass' of the Na+Cl

OpenStudy (anonymous):

which is X grams/moles

OpenStudy (anonymous):

moles = given wt/mol. Wt Mol.wt of nacl=23+35.5=58.5 gm Moles = 29/58.5 = 0.495 = 0.5 moles (approx)

OpenStudy (anonymous):

ok... i got the molar mass as 58.442 so will i do 29 * 58.442

OpenStudy (anonymous):

@aussy123 you've got something a little wrong there hun

OpenStudy (anonymous):

Oh sorry I missed 0.015M